# SCAN: Semantic Document Layout Analysis for Textual and Visual Retrieval-Augmented Generation

### v1

@@ Abstract
With the increasing adoption of Large Language Models (LLMs) and Vision-Language Models (VLMs), rich document analysis technologies for applications like Retrieval-Augmented Generation (RAG) and visual RAG are gaining significant attention. Recent research indicates that using VLMs can achieve better RAG performance, but processing rich documents still remains a challenge since a single page contains large amounts of information. In this paper, we present SCAN (SemantiC Document Layout ANalysis), a novel approach enhancing both textual and visual Retrieval-Augmented Generation (RAG) systems working with visually rich documents. It is a VLM-friendly approach that identifies document components with appropriate semantic granularity, balancing context preservation with processing efficiency. SCAN uses a coarse-grained semantic approach that divides documents into coherent regions covering continuous components. We trained the SCAN model by fine-tuning object detection models with sophisticated annotation datasets. Our experimental results across English and Japanese datasets demonstrate that applying SCAN improves end-to-end textual RAG performance by up to 9.0% and visual RAG performance by up to 6.4%, outperforming conventional approaches and even commercial document processing solutions.
@@
Yuyang Dongthanks: Equal contribution, Nobuhiro Ueda11footnotemark: 1, Krisztián Boros, Daiki Ito, Takuya Sera, Masafumi Oyamada NEC Corporation {dongyuyang, nobuhiro-ueda, krisztian-boros, ito-daiki, takuya-sera, oyamada}@nec.com
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) Lewis et al. (2021); Gao et al. (2024); Fan et al. (2024) technology enables Large Language Models (LLMs) to provide more accurate responses to user queries by retrieving and leveraging relevant knowledge and documents. These knowledge sources such as company financial reports, web pages, insurance manuals, and academic papers often contain complex charts, tables, diagrams, and other non-textual elements, collectively referred to as rich documents. Enabling RAG to accurately support these rich documents is an important research problem.
RAG frameworks are generally categorized into textual RAG and visual RAG. For textual RAG with rich documents, we need to convert documents (PDFs, images, etc.) into text, use a text retrieval, and then generate responses by LLMs. In contrast, visual RAG can directly process rich documents as images using multimodal embedding and retrieval, and then uses Vision Language Models (VLMs) to generate responses.
@@
Figure 2: Overview of applying our SCAN model to current textual and visual RAG systems.
1.1 Challenges
The key part for textual RAG with rich documents is to convert documents into textual formats such as markdown. Recent research Zhang et al. (2025); Fu et al. (2024) indicates that using VLMs for document conversion yields better results than traditional OCR methods. However, it is still a challenge to have a VLM perform text conversion on rich document pages, since a whole page can contain a large amount of complicated content, which can lead to incomplete results and conversion errors.
On the other hand, visual RAG directly retrieves the images of rich document pages, but faces a similar challenge when using a VLM to process visual question answering (VQA) on those rich document pages. This is because user queries often target only specific portions of a document page, meaning that other parts of this page are noise, leading to low accuracy.
Therefore, the common challenge is to use a VLM to process (text conversion or VQA) an entire rich document page containing abundant information at once. One potential solution is to further divide a document page into small regions. Traditional document layout analysis technologies such as DocLayout-YOLO Zhao et al. (2024b) can achieve this objective, but they focus on fine-grained analysis, breaking down content into small components such as titles, paragraphs, tables, figures, and captions. This approach could lose important context when processing isolated components and potentially lead to reduced RAG accuracy. In our experiments, conventional layout analysis methods with VLM text conversion and VQA reduce accuracy by 5.8%–8.7% for textual RAG and 28.7%–40.5% for visual RAG, respectively.
@@ 1.2 Contributions
To address these challenges, we propose SCAN, a novel approach that performs VLM-friendly semantic document layout analysis with “coarse granularity.” Figure 1 compares the layout analysis results between SCAN and a conventional fine-grained method. SCAN can semantically divide regions into boxes that cover continuous components related to the same topic. For example, each of the semantic_box [3], [4], and [5] corresponds to independent topics of IT Services, Social Infrastructure, and Others.
To train a powerful SCAN model, we annotated more than 24k document pages with semantic layout labels. SCAN model is fine-tuned from pre-trained object detection models with this training data. We also designed a specific box selection strategy and post-processing techniques for RAG applications. Figure 2 gives an overview of applying our SCAN model into both textual and visual RAG pipelines. Concretely, each page of an input document is treated as an image and decomposed into semantic chunks by our SCAN model. For textual RAG, the images of semantic chunks are passed to a VLM that performs text conversion, and then input into existing textual RAG systems. For visual RAG, the resulting image chunks can be the retrieval targets that are directly input to existing visual RAG systems.
We test SCAN’s performance using three datasets featuring both English and Japanese documents. Although SCAN is trained on Japanese data, the experiments show that it can achieve good performance on English benchmarks. Experimental results show that in textual RAG, applying SCAN can improve end-to-end performance by 1.7%–9.0%, while in visual RAG, SCAN can enhance end-to-end performance by 3.5%–6.4%. Moreover, although our SCAN requires multiple processes on VLMs rather than a single page-level processing, the total computational time is roughly the same while the total token usage is reduced.
@@ 2.1 Document Layout Analysis
CNN backbones such as DocLayout-YOLO (Zhao et al., 2024b) and transformer backbones such as DiT (Li et al., 2022), LayoutLMv3 (Huang et al., 2022), LayoutDETR (Yu et al., 2023), Beehive (Auer et al., 2024) have been proposed for high-performance document layout analysis which are trained with synthetic and human-annotated datasets Zhong et al. (2019); Li et al. (2020); Pfitzmann et al. (2022). End-to-end document conversion systems have been developed using the above models. Docling (Auer et al., 2024), Marker (Datalab, 2024), and MinerU (Wang et al., 2024) provide a comprehensive pipeline for document layout analysis and text conversions. There are also production systems such as Azure Document Intelligence microsoft (2025) and LlamaParse Premium LlamaParse.AI (2025).
@@ 2.2 VLMs for Document Conversion
Vision-Language Models (VLMs) have emerged as powerful tools for multimodal understanding tasks. Open models such as the Qwen-VL series Bai et al. (2025), or InternVL Chen et al. (2024b) have also demonstrated impressive capabilities in visual document understanding. There are also small-sized OCR-specialized VLMs such as GOT Wei et al. (2024), Nougat Blecher et al. (2023), DocVLM Nacson et al. (2024), and olmOCR Poznanski et al. (2025).
@@ 2.3 Textual and visual RAG
The rapid progress of LLMs has further strengthened RAG, enabling models to inject external knowledge into responses with higher precision and coverage Lewis et al. (2021); Gao et al. (2024); Fan et al. (2024). Typical RAG pipelines in previous works first converted document images or PDFs into plain text, and only then performed indexing and retrieval over the extracted passages Zhang et al. (2025). Recent results show that using VLMs for document text conversion yields better results than traditional OCR tools Zhang et al. (2025); Fu et al. (2024). On the other hand, with the increasing availability of multimodal embedding and VLMs, there is a growing interest in multimodal RAG systems that can directly index and retrieve images, and can leverage VLMs for answer generation Yu et al. (2025); Tanaka et al. (2025); Faysse et al. (2024).
@@ 3 Method
3.1 SCAN Model Training
Given a single page image from a rich document, the semantic document layout analysis problem is to identify and delineate different regions of the image using bounding boxes, each associated with a specific semantic type. This formulation follows classical layout analysis approaches and can be viewed as a specialized object detection task. Accordingly, we fine-tune pre-trained object detection models on a dataset with semantic layout labels, leveraging their robust feature extraction capabilities while adapting them to the requirements of document layout analysis.
3.1.1 Training Data
We define the semantic boxes are regions that contain content unified by a single topic or meaning. Unlike structural divisions (e.g., paragraphs or sections), semantic boxes are defined by their coherent meaning. In other words, all content within a box relates to the same subject and can be understood independently, without needing context from surrounding boxes. A semantic box might contain various structural elements, multiple paragraphs, sections, tables, or charts, as long as they all contribute to the same semantic topic. Besides semantic boxes, we also define global boxes, which correspond to the title, header, footer, date, and author of the document. The global boxes are semantically related to all the elements in the page, while the semantic boxes are more local and can be independent of each other. Figure 1 (right) shows four semantic boxes and three global boxes. Introducing global boxes as well as semantic boxes allows us to represent the page’s semantic dependency structure more accurately.
As we are the first to define semantic document layout analysis in this context, we developed our training and evaluation dataset through a rigorous annotation process. This labor-intensive but crucial step was fundamental to achieving a high-quality model. We collaborated with a specialized data annotation company in Tokyo, engaging six expert Japanese annotators, in a reasonable payment and contract. We asked them to locate the semantic boxes and global boxes in the document images and label them with the corresponding types. The box types we defined are: semantic_box, title, header, footer, date, and author.
@@ 3.1.1 Training Data
Source documents for annotation were collected from Common Crawl PDFs Turski et al. (2023). To ensure diversity in our training data, we first performed clustering on the CC-PDF corpus using image embeddings with MiniCPM-Visual-Embedding Rhapsody Group (2024) model. We then selected a balanced number of candidates from each cluster for annotation, creating a representative dataset that encompasses a wide range of document styles, formats, and domain-specific layouts. We finally labeled 24,577 images and used 22,577, 1,000, and 1,000 samples for training, validation, and test sets. Appendix A provides examples of our dataset.
3.1.2 Object Detection Model Finetuning
There are two popular families of object detection architectures: the CNN-based (YOLO series Redmon et al. (2016); Khanam and Hussain (2024)) and the transformer-based (DETR series Carion et al. (2020); Zhao et al. (2024a)). We tried to fine-tune YOLO11-X (56.9M parameters) [^1] https://docs.ultralytics.com/models/yolo11/ with 5 and RT-DETR-X (67.4M parameters) [^2] https://docs.ultralytics.com/models/rtdetr/ models with our semantic layout annotation data supported on the Ultralytics framework (under New AGPL-3.0 License) to develop our SCAN models.
To evaluate the fine-tuned models in terms of bounding box granularity and precision, we developed a matching-based IoU (Intersection over Union) metric. This metric first uses the Hungarian algorithm to perform bipartite matching between predicted and ground-truth bounding boxes. And then, it calculates an average IoU over the matched bounding box pairs and unmatched bounding boxes. For unmatched bounding boxes, we assigned an IoU of 0 to penalize excessive or insufficient predictions. Table 1 shows the two models’ IoU scores on the validation set of our dataset. Because each predicted bounding box has its confidence score, we varied the confidence threshold to select the optimal predicted bounding box set. With the optimal confidence threshold, the performance of the YOLO11-X and RT-DETR-X finetuned models is similar, while the RT-DETR-X-based model is slightly better. We primarily use RT-DETR-X hereafter.
Model Conf. threshold IoU
YOLO11-X 0.2 53.7
0.3 58.0
0.4 55.8
0.5 49.7
RT-DETR-X 0.2 48.9
0.3 54.8
0.4 58.4
0.5 59.6
@@ 3.1.2 Object Detection Model Finetuning
Table 1: The Intersection over Union (IoU) scores of fine-tuned object detection models. We varied the confidence threshold for predicted bounding boxes from 0.2 to 0.5.
3.2 Predicted Box Selection
Object detection models typically output many bounding boxes along with their corresponding confidence scores. We then select some of the bounding boxes for the final prediction. The simplest selection approach would be to establish a fixed confidence threshold. However, this strategy proves suboptimal for document layout analysis due to document-specific strict constraints:
  • Predicted bounding boxes should cover all textual and visual elements to minimize information loss
  • Predicted bounding boxes should not overlap with each other to minimize information redundancy
This challenge distinguishes document layout analysis from traditional object detection tasks. While conventional object detection aims to identify and localize specific objects with sufficient accuracy, document layout analysis must ensure mutually exclusive and collectively exhaustive predictions.
To address this, we developed an adaptive selection strategy based on two metrics: coverage ratio and non-overlap ratio. The coverage ratio $R_{c}$ measures the proportion of the page area contained within any of the selected boxes:
$R_{c}=\frac{A\left(\bigcup_{i=1}^{n}B_{i}\right)}{A(P)},$ (1)
where $B_{i}$ represents the region of selected box $i$, $A(B)$ represents the area of region $B$, $A(P)$ denotes the whole page area, and $n$ is the number of selected boxes. The non-overlap ratio $R_{o}$ quantifies the degree of redundancy between selected boxes:
$R_{o}=1-\frac{A\left(\bigcup_{1\leq i<j\leq n}(B_{i}\cap B_{j})\right)}{A(P)}.$ (2)
We then define a weighted sum score $S$ to balance these competing objectives:
$S=\alpha\cdot R_{c}+\beta\cdot R_{o},$ (3)
where $\alpha$ and $\beta$ are weighting parameters controlling the trade-off between coverage and overlap.
Our implementation follows a greedy progressive approach: we first establish a minimum confidence threshold to create a candidate pool of boxes. Starting with an empty selection, we iteratively add boxes in descending order of confidence, calculating the score $S$ at each step. A box is added to the final selection if it increases the overall score $S$; otherwise, the box is skipped. This approach ensures we maximize information coverage while minimizing redundancy. We tuned the parameters $\alpha$, $\beta$, and confidence threshold in terms of the IoU score, which resulted in $\alpha=0.9$, $\beta=0.1$, and a confidence threshold of 0.4. We used fine-tuned RT-DETR-X and the validation split of DocSemNet for the parameter tuning.
In addition, to further refine the selection result, we remove boxes completely covered by another selected box, contributing to less redundancy while keeping coverage. After employing these strategies, we achieved an IoU score of 60.2. Appendix C shows the examples before and after the predicted box selection.

### v2

@@ Abstract
With the increasing adoption of Large Language Models (LLMs) and Vision-Language Models (VLMs), rich document analysis technologies for applications like Retrieval-Augmented Generation (RAG) and visual RAG are gaining significant attention. Recent research indicates that using VLMs yields better RAG performance, but processing rich documents remains a challenge since a single page contains large amounts of information. In this paper, we present SCAN (SemantiC Document Layout ANalysis), a novel approach that enhances both textual and visual Retrieval-Augmented Generation (RAG) systems that work with visually rich documents. It is a VLM-friendly approach that identifies document components with appropriate semantic granularity, balancing context preservation with processing efficiency. SCAN uses a coarse-grained semantic approach that divides documents into coherent regions covering contiguous components. We trained the SCAN model by fine-tuning object detection models on an annotated dataset. Our experimental results across English and Japanese datasets demonstrate that applying SCAN improves end-to-end textual RAG performance by up to 9.4 points and visual RAG performance by up to 10.4 points, outperforming conventional approaches and even commercial document processing solutions.
@@
Yuyang Dongthanks: Equal contributionthanks: Current affiliation: SB Intuitions Corp. Email:    yuyang.dong@sbintuitions.co.jp, Nobuhiro Ueda11footnotemark: 1, Krisztián Boros, Daiki Ito, Takuya Sera, Masafumi Oyamada NEC Corporation {dongyuyang,nobuhiro-ueda,krisztian-boros,ito-daiki,takuya-sera,oyamada}@nec.com
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) lewis2021rag; gao2024ragsurvey; fan2024ragllmsurvey technology enables Large Language Models (LLMs) to provide more accurate responses to user queries by retrieving and leveraging relevant knowledge and documents. These knowledge sources, such as company financial reports, web pages, insurance manuals, and academic papers often contain complex charts, tables, diagrams, and other non-textual elements, collectively referred to as rich documents. Effectively enabling RAG systems to understand and utilize such rich, multimodal content remains a key research challenge.
In practice, RAG systems for rich documents follow two major pipeline patterns. Textual RAG first converts documents into text (e.g., Markdown) format, performs text retrieval, and then generates responses with an LLM. Visual RAG, by contrast, retrieves images of pages or regions and uses a Vision-Language Model (VLM) to read the images and generate an answer directly. Although their modalities differ, both pipelines ultimately depend on VLMs—either to convert visual regions into text or to interpret them directly—and both tend to break down when an entire page is processed at once.
@@
Figure 2: Overview of applying our SCAN model to current textual and visual RAG pipelines.
Therefore, a common challenge is having a VLM process an entire rich document page (text conversion or VQA) at once. One potential solution is to further divide a document page into small regions. Traditional document layout analysis technologies such as DocLayout-YOLO doclayoutyolo can achieve this objective, but they focus on fine-grained analysis, breaking down content into small components such as titles, paragraphs, tables, figures, and captions. This approach could lose important context when processing isolated components and potentially lead to reduced RAG accuracy. In our experiments, conventional layout analysis methods with VLM text conversion and VQA degraded the RAG performance in most cases.
@@ 1.1 Contributions
To address these challenges, we propose SCAN, a novel approach that performs VLM-friendly semantic document layout analysis with “coarse granularity.” Figure 1 compares the result of conventional layout analysis with that of SCAN. SCAN can semantically divide regions into boxes that cover contiguous components related to the same topic. For example, each of the semantic box [3], [4], and [5] corresponds to independent topics of IT Services, Social Infrastructure, and Others.
To train a powerful SCAN model, we annotated more than 24k document pages with semantic layout labels. SCAN model is fine-tuned from pre-trained object detection models with this training data. We also designed post-processing techniques for RAG applications. Figure 2 gives an overview of applying our SCAN model to both textual and visual RAG pipelines. Concretely, each page of an input document is treated as an image and decomposed into semantic chunks by our SCAN model. For textual RAG, the images of semantic chunks are passed to a VLM that performs text conversion, and then input into existing textual RAG systems. For visual RAG, the resulting image chunks can be the retrieval targets that are directly input to existing visual RAG systems.
We evaluate SCAN’s performance using three datasets featuring both English and Japanese documents. Although SCAN is trained on Japanese data, the experiments show that it can achieve good performance on English benchmarks. Experimental results show that in textual RAG, applying SCAN can improve end-to-end performance by 2.7–9.4 points, while in visual RAG, SCAN can enhance end-to-end performance by 5.6–10.4 points. Moreover, although our SCAN requires multiple inferences on VLMs rather than single page-level processing, the total computational time is reduced because of the reduced number of input tokens per inference.
@@ 2.1 Document Layout Analysis
CNN-based models such as DocLayout‐YOLO (doclayoutyolo) and Transformer-based models such as DiT (li2022dit), LayoutLMv3 (huang2022layoutlmv3), and Beehive (docling) have been proposed for high-performance document layout analysis which are trained with synthetic and human-annotated datasets publaynet; docbank; doclaynet2022. Building on these models, several end-to-end document conversion systems have been developed. Docling (docling), Marker (marker), and MinerU (mineru) provide a comprehensive pipeline for document layout analysis and text conversion. In addition, production systems such as Azure Document Intelligence adi and LlamaParse Premium llamaparse are available.
@@ 2.2 VLMs for Document Conversion
Vision-Language Models (VLMs) have emerged as powerful tools for multimodal understanding tasks. Open models such as the Qwen-VL series Qwen2.5-VL; qwen3-vl or InternVL internvl have also demonstrated impressive capabilities in visual document understanding. Moreover, smaller OCR-specialized VLMs, including GOT got, Nougat nougat, DocVLM docvlm, and olmOCR olmocr, have been developed to efficiently handle document text extraction.
@@ 2.3 Textual and Visual RAG
The rapid progress of LLMs has further strengthened RAG, enabling models to inject external knowledge into responses with higher precision and coverage lewis2021rag; gao2024ragsurvey; fan2024ragllmsurvey. Typical RAG pipelines in previous works first converted document images or PDFs into plain text, and only then performed indexing and retrieval over the extracted passages ohr. Recent results show that using VLMs for document text conversion yields better results than traditional OCR tools ohr; orcbenchv2. On the other hand, with the increasing availability of multimodal embedding and VLMs, there is a growing interest in multimodal RAG systems that can directly index and retrieve images, and can leverage VLMs for answer generation visrag; vdocrag; colpali.
@@ 3 Method
Our goal is to design a VLM-friendly layout analysis module that divides a rich document page into a small number of semantically coherent regions with coarse granularity. We call this task semantic document layout analysis, and the resulting regions semantic chunks. In contrast to conventional layout analysis, which produces many fine-grained boxes for structural elements (titles, paragraphs, tables, figures, etc.), SCAN aims at a coarser granularity that better matches how humans understand a page and how RAG systems use it.
A semantic chunk on a page is defined as a region whose content is unified by a single subtopic. Complex pages often contain several subtopics under one broader topic. For example, in the page shown in Figure 1, the overall topic is “business areas of a company,” but there are three distinct subtopics: IT Services, Social Infrastructure, and Others. Each of these subtopics is represented by one semantic chunk. Formally, we require that everything inside one chunk is necessary to understand that subtopic and that no important part of the subtopic lies outside the chunk. This notion is different from structural divisions such as paragraphs, sections, or table cells. Structural divisions are defined by superficial document elements (e.g., line breaks or section headers) and do not necessarily align with topical boundaries. They are particularly unreliable for floating elements such as figures and tables, or for infographic-style pages where text and graphics are freely laid out. Therefore, we define semantic chunks directly in terms of topical coherence rather than layout structure. In our implementation, semantic chunks are represented as rectangular bounding boxes, and we introduce two types of boxes. Semantic boxes correspond to content related to one subtopic as defined above; one page typically contains several semantic boxes. Global boxes correspond to page-level metadata such as the title, header, footer, date, and author. Global boxes are semantically related to the whole page, whereas semantic boxes are local and mostly independent of each other. Figure 1 (right) illustrates five semantic boxes and three global boxes produced by our SCAN model. Introducing global boxes as well as semantic boxes allows us to represent the page’s semantic dependency structure more accurately.
Given a single-page image from a rich document, our semantic document layout analysis task is to predict a set of such bounding boxes together with their box types. This setup follows classical layout analysis and can be formulated as a multi-class object detection problem. Accordingly, we fine-tune pre-trained object detection models on a dataset with semantic layout labels. In the rest of this section, we first describe how we construct the dataset used to train SCAN (Section 3.1), then explain how we fine-tune object detection models (Section 3.2), and finally detail the post-processing applied when integrating SCAN into textual and visual RAG pipelines (Section 3.3).
Cluster Domain # Pages # Boxes
Tables, Charts 3,766 17,024
Flyers, Magazines, Menus, Recipes 3,747 21,555
Maps, Traveling information 3,383 16,558
Itemized documents 3,771 19,062
Handwritten text 821 3,481
Vertical texts 3,768 16,520
Math 1,213 5,479
Manuals, Guidelines, Blueprints 4,108 23,472
Table 1: Statistics of document domains in our dataset.
Box Type Avg. Boxes Per Page # Boxes
semantic box 3.13 76,862
title 0.49 12,137
header 0.38 9,416
footer 0.73 17,969
date 0.12 2,888
author 0.16 3,879
All types 5.01 123,151
Table 2: Statistics of box types in our dataset.
3.1 Dataset Construction
@@
As we are the first to define semantic document layout analysis in this context, we developed our training and evaluation dataset through a rigorous annotation process. We first collected document pages from the Japanese portion of Common Crawl PDFs ccpdf. To ensure diversity in our training data, we first performed clustering on the CC-PDF-JP corpus using image embeddings with the MiniCPM-Visual-Embedding minicpmv model. We then selected balanced number of sample pages from each cluster and included them in our dataset; the statistics for each cluster are provided in Table 1. We finally labeled 24,577 pages and split them into 22,577 pages for training, 1,000 for validation, and 1,000 for testing. Appendix A provides examples of our dataset.
The annotation task required annotators to identify semantic chunks within each page and draw bounding boxes around them, followed by assigning a box type. The box types we defined are semantic box, title, header, footer, date, and author. Table 2 reports the number of boxes for each class and the average number of boxes per page.
We collaborated with a specialized data annotation company in Tokyo, engaging six expert Japanese annotators, with a reasonable payment and contract. We first conducted a pilot annotation on small number of samples to establish a detailed guideline regarding box granularity, segmentation criteria, and the handling of ambiguous cases. During the pilot stage, we measured inter-annotator agreement using a matching-based IoU (Intersection over Union) metric (described in Section 3.2) and obtained a score above 0.7. Because the IoU metric is relatively sensitive to box misalignment, we considered the agreement sufficiently high. Thus, we adopted a single-annotator-per-sample protocol for the main annotation phase. We also regularly checked annotation quality and updated the guideline when unclear. Appendix B provides details of our annotation guideline.
Model Confidence IoU (%) Coverage (%)
YOLO11-X 0.2 53.7 81.0
0.3 58.0 79.0
0.4 55.8 73.1
0.5 49.7 64.0
RT-DETR-X 0.2 48.9 78.0
0.3 54.8 78.8
0.4 58.4 79.0
0.5 59.6 77.7
@@
Table 3: The Intersection over Union (IoU) scores and coverage ratios of fine-tuned object detection models. We varied the confidence threshold for predicted bounding boxes from 0.2 to 0.5.
Dataset Domain Lang. # Docs # Pages # QA QA Type
OHR-Bench
Academic, Administration, Finance,
Law, Manual, Newspaper, Textbook
En 1,261 8,561 8,498
Text, Table, Formula,
Chart, Reading Order
Allganize
Finance, Information Technology, Manufacturing,
Public Sector, Retail/Distribution
Ja 64 2,167 300 Text, Chart, Table
BizMMRAG
Finance, Government, Medical,
Consulting Sectors, Wikipedia
Ja 42 3,924 146 Text, Chart, Table
Table 4: Statistics of our evaluation datasets.
3.2 Object Detection Model Fine-tuning
There are two popular families of object detection architectures: the CNN-based (YOLO series yolo1; yolo11) and the Transformer-based (DETR series detr-eccv-2020; rtdetr). We fine-tuned YOLO11-X (56.9M parameters) [^1] https://docs.ultralytics.com/models/yolo11/ and RT-DETR-X (67.4M parameters) [^2] https://docs.ultralytics.com/models/rtdetr/ supported in the Ultralytics framework (under New AGPL-3.0 License) to develop our SCAN models.
To evaluate the fine-tuned models in terms of bounding box granularity and precision, we developed a matching-based IoU (Intersection over Union) metric. This metric first uses the Hungarian algorithm to perform bipartite matching between predicted and ground-truth bounding boxes. And then, it calculates an average IoU over the matched bounding box pairs and unmatched bounding boxes. For unmatched bounding boxes, we assigned an IoU of 0 to penalize excessive or insufficient predictions.
Table 3 shows the two models’ IoU scores on the validation set of our dataset. Because each predicted bounding box has its confidence score, we varied the confidence threshold to select the optimal predicted bounding box set. We also computed the coverage score of the selected bounding boxes, defined as the ratio of the area covered by the selected boxes to the total area of all ground-truth boxes. With the optimal confidence threshold, the IoU score of the RT-DETR-X fine-tuned models is better than that of the YOLO11-X fine-tuned models. RT-DETR-X (confidence: 0.5) achieved the best IoU score of 59.6. However, its coverage score is the lowest (77.7) among all the RT-DETR-X settings. Thus, we selected the model based on RT-DETR-X with a confidence threshold of 0.4 as our primary SCAN model. Appendix D shows examples of the predicted bounding boxes by our SCAN model.